OpenStudy (anonymous):

solve square root of28 by division

OpenStudy (anonymous):

i'm not sure what you mean by division but you can surely estimate: \(\sqrt{28}\) is between \(\sqrt{25}\) and \(\sqrt{36}\) so that means that the root has to be between \(5\) and \(6\) and since \(28\) is closer to \(25\) then it is to ](36\) the answer can be narrowed down to being between \(5\) and \(5.5\), maybe even less now if you use a calculator, you get the answer: \(\sqrt{28}\approx5.3\)

OpenStudy (caozeyuan):

sqrt(28)=sqrt(7*4)=sqrt(7)*sqrt(4)+2*sqrt(7) is that what you need?

OpenStudy (caozeyuan):

oh! a typo made! it should be =2*sqrt(7), not + 2*sqrt(7)